Exercise 6-4

Comfi Airways, Inc., a small two-plane passenger airline, has asked for your assistance in some basic analysis of its operations. Both planes seat 10 passengers each, and they fly commuters from Comfi’s base airport to the major city in the state, Metropolis. Each month 40 round-trip flights are made. Shown below is a recent month’s activity in the form of a cost-volume-profit income statement.

Fare revenues (400 fares) $47,400

Variable costs

Fuel 15,614

Snacks and drinks 740

Landing fees 2040

Supplies and forms 1040 19434

Contribution margin 27966

Fixed costs

Depreciation 3090

Salaries 15109

Advertising 400

Airport hanger fees 1520 20119

Net income 7847

(a) Calculate the break-even point in (1) dollars and (2) number of fares. (Round answers to 0 decimal place, e.g. 1,225.)

1. break-even point $________?

2. break-even point __________(fares)?

(b) Without calculations, determine the contribution margin at the break-even point.

1. break even point$_______?

(c) If fares were decreased by 10%, an additional 100 fares could be generated. However, total variable costs would increase by 20%. (Round answers to 0 decimal place, e.g. 1,225.)

(1) How much would net income be impacted by this change?

Net income (increases or decreases?) ______ to $ ____________?
Answer #1

1) Contribution margin ratio = 27966/47400 = 59%

Break even sales = Fixed cost/Contribution margin ratio = 20119/.59 = $34100

Break even unit = 34100/118.5 = 288 fares

2) Contribution margin at break even = Fixed cost = 20119

3) Fare = 118.5*90% = 106.65

Variable cost = 48.585*1.2 = 58.302

Net income = (500*48.348)-20119 = 4055

Net income decrease to $4055